OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

\[r^2s^4t/r^2st^3\]

OpenStudy (anonymous):

Uh...

OpenStudy (anonymous):

dividing monomials

OpenStudy (anonymous):

Yeah i dont know what yet :( sorry

OpenStudy (anonymous):

thats ok!

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

You mean \[\frac{r^2s^4t}{r^2st^3}\] right?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Do you know about the properties of exponents with a common base in division? \[\frac{x^n}{x^m} = x^{n-m}\] You can rewrite your problem as \[\frac{r^2}{r^2} * \frac{s^4}{s} * \frac{t}{t^3}\]Use the property I just mentioned to simplify each fraction, then multiply the results together for your answer. Remember also that \[x^{-n} = \frac{1}{x^n}\]and\[\frac{1}{x^{-n}} = x^n\]

OpenStudy (anonymous):

yup, im just confused on how i am supposed to simplify it.

OpenStudy (anonymous):

is it  ?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@jdoe0001

OpenStudy (whpalmer4):

Yes, that is correct. You could also write it as \[s^3t^{-2}\]if you prefer to have it all on one line.

OpenStudy (anonymous):

i solved this one is it right? \[3x^6 / 18x\]

OpenStudy (whpalmer4):

Is that the problem or the answer?

OpenStudy (anonymous):

\[x^5/6\]

OpenStudy (anonymous):

is that the correct answer?

OpenStudy (whpalmer4):

Yes, that's correct. Sorry, didn't see the notification that you'd responded.